OpenStudy (karisleft):

Evaluate the expression under the given conditions. sin(θ − ϕ); tan θ=12/5, θ in quadrent III, sin ϕ = -sqrt10/10, ϕ in quadrent IV

OpenStudy (karisleft):



OpenStudy (karisleft):

looks like that if it helps

OpenStudy (shadowlegendx):

@ParthKohli

Parth (parthkohli):

yeah you just need to calculate \(\cos \theta, \sin \theta, \cos \phi, \sin \phi\) but that shouldn't be troublesome because you're given all the data you need

Parth (parthkohli):

the quadrants tell you the sign of the trigonometric ratios in a way, since a quadrant will uniquely determine the sign of your trig functions. and whenever you're given tan/cot, you can make a triangle to calculate the magnitudes of sine and cosine. i'd speak more on this, but tell me whether you've encountered this before

OpenStudy (karisleft):

this is the first, its a practice question in my book, trying to understand

Parth (parthkohli):

If this is the first problem in your practice book, then you're using the wrong practice book.

OpenStudy (karisleft):

jeeez, well where do i start

satellite73 (satellite73):

use the subtraction angle formula \[\sin(A-B)=\sin(A)\cos(B)-\sin(B)\cos(A)\]

satellite73 (satellite73):

you got \(\tan(\theta)=\frac{12}{5}\) from that you should be able to find \(\cos(\theta)\) and \(\sin(\theta)\)